Subject: APO 3.5m users committee minutes, June 2010
From: Suzanne Hawley
Submitted: Sat, 19 Jun 2010 12:16:06 -0700 (PDT)
Message number: 1188
(previous: 1187,
next: 1189
up: Index)
APO 3.5-m Users Committee Phonecon, 6/14/10
Attending: Suzanne Hawley, Mark Klaene, Michael Strauss, Bill Ketzeback,
Jon Holtzman, Remy Indebetouw, Scott Anderson, Bruce Gillespie
Absent: Al Harper, John Bally, Jon Fulbright
**********************************
User feedback, comments from institutional representatives:
Chicago (Al Harper) - no report.
New Mexico State (Jon Holtzman) - Jon had nothing new to report.
Washington (Scott Anderson) - Scott had nothing new to report.
Colorado (John Bally) - no report.
Princeton (Michael Strauss) - Michael said "we've been happy."
Johns Hopkins (Jon Fulbright) - no report.
Virginia (Remy Indebetouw) - Remy mentioned that a UVa user has been
repeatedly "clouded out" during scheduled observing time. Suzanne and
Remy said that they are trying to compensate by using DD time and raising
the internal priority at Virginia.
**********************************
Discussion of telescope/instruments report - Klaene:
Mark highlighted the recent cloudy weather and the return of Miller
moths -- whether the new drives are better or worse when they run over
a moth has yet to been determined. The main issue this past month has
been azimuth axis drive resonances during slews and drive startup. A low-pass
filter was installed in the drive system, and after adding an additional
amplifier the drive system is working much better. The altitude axis
also experiences resonances that are excited at low altitudes; it will get
the same treatment as the azimuth axis. For the instruments, there were no
significant problems, although we saw instrument temperature fluctuations due
to the warm ambient air temperatures. Given the recent and long-term problems
we have had with mechanical cryo-coolers, Suzanne and Mark stated that
we will not consider such cooling systems for future instruments. The
NIC-FPS fan-out board is still being worked on at Colorado. The spare telescope
drive box and motor will be ready in July. Improving high-altitude tracking
remains in the work plan.
*************************
3.5-m Telescope, Instruments, and CIF Projects Highlights, 5/4/10 through 6/10/10
0) Overview
Spring finally showed up bringing a heavy pollen season and lots of Miller
moths. Wind has generally been milder than some seasons and we have
actually seen a little rain. Blowing sand has also generally been lower
than a typical spring. After a lot of work by Fritz and others we seem
to have made major strides in improving the axis controller performance
on the azimuth axis. The problems we were seeing did not affect a lot of
users but were certainly irritating and needed correction.
1) Telescope
Fritz managed to get the low-pass filter in the az controller servo loop
but it was attenuating the servo drive signal too much so an external
amplifier had to be installed to provide additional gain for the
post-filtered signal. This allowed for better servo tuning and resulted
in reduced tracking error and resonance. A similar approach is now
planned for altitude.
We are still investigating improvements in high altitude tracking but
little progress has been made due to weather and other priority issues.
2) Instruments
SPIcam, DIS, TripleSpec, NIC-FPS, the Echelle, and Agile ran with no
serious problems. Due to some very hot daytime temperatures DIS and
SPIcam and even Echelle temperatures were running higher than we like
during the day and only after sunset and dome ventilation recovered to
acceptable limits. We have been unable to find anything wrong with the
camera cooling systems but are placing instruments on the vacuum pump
when available to try and improve these temperature transients.
The new NIC-FPS fanout board testing at CU has been slowed due to problems
with the spare Leach controller. The controller was sent back to Bob Leach
who found a couple of problems. The output is now stable but there is a
discrepancy in the gains between the two channels that is still being
investigated.
3) CIF projects
The spare telescope drive box machining and fabrication continues; some
parts are assembled and additional spare components have been delivered.
A complete spare motor assembly will be finished this summer.
We have an RFP out to upgrade the 5-ton hoist in the dome that is used
for M1 mirror movements to a variable frequency drive. The new drive
will give us finer control when removing and re-installing the primary mirror.
High-altitude tracking is still being investigated. We have confirmed
that M1 comes off the transverse actuators as early as 82 degrees and
that we have seen 30 or more arcsec of image motion that the
auto-guider cannot correct when operating above 85 deg altitude.
**********************************
Summer engineering update - Klaene:
Mark said that the summer shutdown will be conducted from 26 July
through ~14 August (weather permitting). This is a relatively short shutdown,
which will entail washing the primary mirror, servicing and recoating
the secondary and tertiary mirrors, and annual instrument maintenance.
Since we're coming out of the shutdown during the monsoon season, the
date that we return to science observing will depend strongly on the weather.
**********************************
Echelle upgrade update - Gillespie:
We are tracking down the echelle documentation so that we have a
baseline collection before launching into the upgrade. Yerkes has just
sent us a couple of boxes of documents, and we have identified other
echelle materials currently in the custody of Shu-I Wang, Lew Hobbs, Don
York, and Dan Schroeder. All the documentation should be on-site
at APO by the end of summer. Jon Holtzman is leading the echelle upgrade
project, and is preparing a description of the upgrade plan. A
preliminary design review of the upgrade is planned for this fall.
**********************************
APO futures committee - Hawley, Gillespie:
An ARC Futures Team, consisting of Mike Shull, Suzanne, Daniel Eisenstein,
Michael Strauss, and Bruce have begun the process of determining APO's
future in the post-2014 time frame. The main elements being considered
are soliciting a program for the 2.5-m telescope after SDSS-III, the
evolution of the ARC institutions and shares, and new instrumentation.
A larger Futures Committee needs to be formed, with representatives from
the ARC member institutions plus a subset of the SDSS-III collaboration.
This committee will be tasked with reviewing the Team's recommendations,
and providing an information conduit back and forth with the 3.5-m and
SDSS-III constituencies.
**********************************
Summer meeting schedule - Hawley:
As is our tradition to skip one of the monthly meetings in the summer,
we will skip July and hold our next Users Committee telecon on 16 August,
after the summer shutdown activities are completed.
**********************************
Miscellaneous topics:
Bruce mentioned that Bill Ketzeback now has the long-lost Gascoyne
corrector lenses that were originally manufactured (back in the late-80s?),
which could enable the 3.5-m to produce a fairly wide-field corrected
focal plane. We don't know much about the optical properties of the
lenses, but these could be reverse-engineered.
Remy asked if anyone is using TripleSpec in extended-source mode, via
scripts. Michael Strauss said that the Princeton users would be very
interested in this capability. Jon Holtzman thought it would be a great
thing to have, and could be implemented by borrowing from NIC-FPS software.
Suzanne said that we should look into this as a high priority.
**********************************
ACTION ITEMS:
[open from previous meeting]:
===> ACTION: Users Committee members should poll their users to find
out how important it is to their science programs to track targets within
5 degrees of zenith.
STATUS: Comments were heard from three institutions, which basically
said it would be nice, but not a big deal, to go to higher elevations.
The action will remain open until next meeting for additional user comments.
===> ACTION: Users Committee members should forward the NIIS white paper
to their users, and potential users should send any comments or questions
to Suzanne Hawley.
STATUS: This action is CLOSED. No user comments were received.
When it is ready, NIIS will come to APO as a visiting instrument.
[new actions from this meeting]:
None
**********************************
Next meeting: The next Users Committee phonecon will be on Monday 16
August at 8:30 AM Pacific Time. The agenda and other materials will be
sent to the committee members during the preceding week.
APO APO APO APO APO Apache Point Observatory 3.5m APO APO APO
APO
APO This is message 1188 in the apo35-general archive. You can find
APO the archive on http://www.astro.princeton.edu/APO/apo35-general/INDEX.html
APO To join/leave the list, send mail to apo35-request@astro.princeton.edu
APO To post a message, mail it to apo35-general@astro.princeton.edu
APO
APO APO APO APO APO APO APO APO APO APO APO APO APO APO APO APO